OpenStudy (anonymous):

PLEASE help me set this problem up A pyramid has a rectangular base with a length that is 2 m more than its width. The height of the pyramid is 6m, and its volume is 48 meters cube. Find the length and width

OpenStudy (cwrw238):

Volume of a pyramid = (1/3) * height* area of the base

OpenStudy (cwrw238):

let the width be x m then the length = (x+ 2) m form an equation using the formula for the volume and solve for x.

OpenStudy (anonymous):

thanks, so the equation will be 48=1/3(6)x(x+2)? or am i wrong?

OpenStudy (cwrw238):

that is correct

OpenStudy (cwrw238):

this will lead to a quadratic equation

OpenStudy (anonymous):

okay thanks, do i have to multiply 6 to x or leave it on it's own ?

OpenStudy (cwrw238):

(1/3)*6 = 2

OpenStudy (anonymous):

okay thank you so much!!!!

OpenStudy (cwrw238):

when you are writing a fraction all in one line put it in parentheses to avoid ambiguity

OpenStudy (anonymous):

so once i divide six from 1/3 and get 2 do i multiply it to x(x+2)?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thank you so much once again (:

OpenStudy (cwrw238):

yw you'll get a quadratic equation with 2 roots Ignore the negative root. x will be the width of the base and length will be x+2